# Practical learning of multi-time statistics in open quantum systems

### v1

@@ 5.2 Dynamically Sampling Non-Markovian Open Quantum Systems
We perform a demonstration that consolidates our tools and reconstructs a long non-Markovian process simulated on a quantum device. Specifically, we consider one end of a Heisenberg-coupled spin-chain and estimate the 20-step/42-body process tensor Choi state using the finitely correlated state ansatz [57]. This method reconstructs the 1D global quantum states, within the finitely correlated approximation, using only the marginals of an odd number neighbouring spins that satisfy certain left and right invertibility conditions. As an ansatz, to describe an $n$-body state, choose left and right ranks to be $d^{2l}$ and $d^{2r}$ respectively. Here, $d$ is dimension each local body. This requires $R$-body marginals with $R=l+r+1$, leading to a total of $n-R+1$ local marginals. In the process tensor context, we start with $k$ steps process, represented by a $12k+$ 1-body state. If $l=r:=\ell$, then the problem requires estimation of all $\ell$ step marginals: $\{\Upsilon_{\ell:0}$, $\Upsilon_{\ell+1:1}$, $\cdots$, $\Upsilon_{k:k-\ell}\}$. With the input of classical shadows, these $k-\ell+1$ marginals may be simultaneously estimated, leaving the characterisation scaling logarithmic in the number of steps.
Full details of the theorem reconstructing a state from its marginals may be found in Ref. [57], however we reconfigure the statements here in the form of a graphical diagram in Figure 5 to reconstruct finitely correlated process ansatz. The procedure involves computing pseudoinverses to connect together process tensor marginals. This further motivates the need for a robust marginal estimate, since inverses can only be computed with complete tomographic information about a state. We start with $\Upsilon_{\ell:0}$ and move one leg across to obtain the next $12\ell+$ 1 body marginal $\text{Tr}_{\mathfrak{o}_{\ell+1},\mathfrak{o}_{0}}[\Upsilon_{\ell+1:0}]$. However, causality constraints mean that this term reduces to $\mathbb{I}_{\mathfrak{i}_{\ell+1}}\otimes\text{Tr}_{\mathfrak{o}_{0}}[\Upsilon_{\ell:0}]$, which is information we already have. Moving one leg over, the next marginal is $\Upsilon_{\ell+1:1}$. This continues until we get to $\Upsilon_{k:k-\ell}$. Collectively, these form the set
@@
The connections of these terms to construct the overall finitely correlated process is depicted in Figure 5a, with the effective many-time state depicted in Figure 5b. The net result is a matrix product operator (MPO) of the form
$\Upsilon^{(fcp)}_{k:0}=\sum_{\lambda_{0},\ldots,\lambda_{k-1}}\Gamma_{0}^{(\lambda_{0})}\otimes\Gamma_{1:0}^{(\lambda_{0},\lambda_{1})}\otimes\cdots\otimes\Gamma_{k-1:k-2}^{(\lambda_{k-2},\lambda_{k-1})}\otimes\Gamma_{k:k-1}^{(\lambda_{k-1})}.$ (41)
Matrices $\Gamma$s represent quantum channels between neighbouring steps. The $\lambda$s represent the bond of the MPO and carry temporal correlations.
@@ 5.3 Reconstructing finitely correlated process ansatz on NISQ devices
To validate that the estimate does indeed accurately capture environmental dynamics, 170 different sequences of random instruments are then run on the device at 16 3841638416\>38416 384 shots. The outcomes of the instruments across the 21 different times are collected and used to generate different marginal probability distributions. The effect of marginalising over the outcomes at a given time is that the operation applied is a deterministic CPTP channel. The marginal probability distributions for each sequence of operations is predicted by our process characterisation and compared via Hellinger fidelity to the actual results.
Figure 7: a We characterise a 20-step process of a non-Markovian quantum environment on ibm_lagos and construct its finitely correlated state representation. b This characterisation is then validated against 170 sequences of random instruments: the reconstructed process tensor is used to predict different $l$-time marginal probability distributions of the instrument outcomes and compared against the experimental results. c Properties of the system when left idle as a function of time step, showcasing conventional measures of non-Markovianity. We observe the non-monotonic behaviour of both the purity of the system, and the trace distance between two states initialised as $Z$ and $Y$ eigenstates. d Properties of the 3-step marginal process tensors as a function of time, we measure both the total memory and the temporal quantum entanglement, demonstrating a genuinely complex non-Markovian process.
@@
for the measurement outcomes at the different times. We first sample from this distribution for a given $\mathbf{J}_{20:0}$ by running the sequence on the device at 16 3841638416\>38416 384 shots and recording each of the results. Naturally, at $2\times 10^{6}$ outcomes, this distribution is already too large to sensibly measure and compare. Hence, we consider marginals of the distribution. A marginal of length $\ell$ at a leading time $j$ is the marginal distribution of Eq. (42)
@@
which, for the resulting graphical diagram in Figure 5, can be contracted efficiently. Note that where we have dropped the instrument superscript, this indicates that we have discarded the ancilla outcome, effecting a CPTP map: $\mathcal{J}_{i}=\mathcal{J}_{i}^{(0)}+\mathcal{J}_{i}^{(1)}$. The boxplots of Figure 7b display the Hellinger distance between the quantum computer sample, and Eq. (44) for each of the 170 sequences, for each marginal $1-5$, and for each time $0-20$.

### v2

@@ 5.2 Dynamically Sampling Non-Markovian Open Quantum Systems
We perform a demonstration that consolidates our tools and reconstructs a long non-Markovian process simulated on a quantum device. Specifically, we consider one end of a Heisenberg-coupled spin-chain and estimate the 20-step/42-body process tensor Choi state using the finitely correlated state ansatz [59]. This method reconstructs the global quantum state, within the finitely correlated approximation (i.e., that the correlation structure is one dimensional and exponentially decreasing with distance), using only the marginals of an odd number of neighbouring spins that satisfy certain left and right invertibility conditions. More specifically, this compact representation is a subclass of matrix product operators (MPOs). Further information on MPOs and other tensor network states may be found in Ref. [60]. As an ansatz, to describe an $n$-body state, choose left and right ranks to be $d^{2l}$ and $d^{2r}$ respectively (where left and right ranks are the dimensions of the image of a matrix acting as a map to the left or right, respectively). Here, $d$ is the dimension of each local body. This requires $R$-body marginals with $R=l+r+1$, leading to a total of nR+1n-R+1 local marginals. In the process tensor context, we start with $k$-step process, represented by a $2k+1$-body state. If $l=r:=\ell$, then the problem requires estimation of all $\ell$ step marginals: $\{\Upsilon_{\ell:0}$, $\Upsilon_{\ell+1:1}$, $\cdots$, $\Upsilon_{k:k-\ell}\}$. With the input of classical shadows, these $k-\ell+1$ marginals may be simultaneously estimated, leaving the characterisation scaling logarithmically in the number of steps.
Full details of the theorem reconstructing a state from its marginals may be found in Ref. [59], however, we reconfigure the statements here in the form of a graphical diagram in Figure 5 to reconstruct a finitely correlated process ansatz. The procedure involves computing pseudoinverses to connect together process tensor marginals. This further motivates the need for a robust marginal estimate, since inverses can only be computed with complete tomographic information about a state. A description of the Ref. [59] procedure as adapted to process tensors is as follows. We start with $\Upsilon_{\ell:0}$ and move one leg across to obtain the next $2\ell+1$ body marginal $\text{Tr}_{\mathfrak{o}_{\ell+1},\mathfrak{o}_{0}}[\Upsilon_{\ell+1:0}]$. However, causality constraints mean that this term reduces to $\mathbb{I}_{\mathfrak{i}_{\ell+1}}\otimes\text{Tr}_{\mathfrak{o}_{0}}[\Upsilon_{\ell:0}]$, which is information we already have. Moving one leg over, the next marginal is $\Upsilon_{\ell+1:1}$. This continues until we get to $\Upsilon_{k:k-\ell}$. Collectively, these form the set
@@
The connections of these terms to construct the overall finitely correlated process is depicted in Figure 5a, with the effective many-time state depicted in Figure 5b. The net result is an MPO of the form
$\Upsilon^{(fcp)}_{k:0}=\sum_{\lambda_{0},\ldots,\lambda_{k-1}}\Gamma_{0}^{(\lambda_{0})}\Gamma_{1:0}^{(\lambda_{0},\lambda_{1})}\cdots\Gamma_{k-1:k-2}^{(\lambda_{k-2},\lambda_{k-1})}\Gamma_{k:k-1}^{(\lambda_{k-1})}.$ (41)
[…]
@@ 5.3 Reconstructing finitely correlated process ansatz on NISQ devices
To validate that the estimate does indeed accurately capture environmental dynamics, 170 different sequences of random instruments are then run on the device at 16 38416\>384 shots. The outcomes of the instruments across the 21 different times are collected and used to generate different marginal probability distributions. The effect of marginalising over the outcomes at a given time is that the operation applied is a deterministic CPTP channel. The marginal probability distributions for each sequence of operations is predicted by our process characterisation and compared via Hellinger fidelity to the actual results.
Figure 7: a We characterise a 20-step process of a non-Markovian quantum environment on ibm_lagos and construct its finitely correlated state representation, in Eq. (41). b This characterisation is then validated against 170 sequences of random instruments: the reconstructed process tensor is used to predict different $l$-time marginal probability distributions of the instrument outcomes and compared against the experimental results. c Properties of the system when left idle as a function of time step, showcasing conventional measures of non-Markovianity. We observe the non-monotonic behaviour of both the purity of the system, and the trace distance between two states initialised as $Z$ and $Y$ eigenstates. d Properties of the 3-step marginal process tensors as a function of time, we measure both the total memory and the temporal quantum entanglement, demonstrating a genuinely complex non-Markovian process.
@@
for the measurement outcomes at the different times. We first sample from this distribution for a given $\mathbf{J}_{20:0}$ by running the sequence on the device at 16 38416\>384 shots and recording each of the results. Naturally, at $2\times 10^{6}$ outcomes, this distribution is already too large to sensibly measure and compare. Hence, we consider marginals of the distribution. A marginal of length $\ell$ at a leading time $j$ is the marginal distribution of Eq. (42)
@@
which, for the resulting graphical diagram in Figure 5, can be contracted efficiently. Note that where we have dropped the instrument superscript, this indicates that we have discarded the ancilla outcome, effecting a CPTP map: $\mathcal{J}_{i}=\mathcal{J}_{i}^{(0)}+\mathcal{J}_{i}^{(1)}$. The boxplots of Figure 7b display the Hellinger distance between the quantum computer sample, and Eq. (44) for each of the 170 sequences, for each marginal 151-5, and for each time 0200-20.